OpenStudy (anonymous):

the sum of two consecutive even integers is -298 what are the integers

OpenStudy (anonymous):

x + x + 2 = -298 (now solve for x)

OpenStudy (anonymous):

2x+2=-298 2x=-300 x=-150

OpenStudy (anonymous):

I think I got the answer is it -150, -60, -88

OpenStudy (anonymous):

your two integers are x and x + 2. you already figured x to be -150. So the other integer, x + 2, is -150 + 2 = -148. check... -150 + (-148) = -298

OpenStudy (anonymous):

its is the right answer thx for ur help

OpenStudy (anonymous):

anytime :)